A perfect gas has a molar mass of 26 kg/kmol and a value of 𝛾 = 1.26. Calculate the heat rejected when unit mass of the gas is contained in a rigid vessel at 3 bar and 315℃, and is then cooled until the pressure falls to 1.5 bar. Assuming that 𝐢𝑣 = 1.229 π‘˜π½/π‘˜π‘”πΎ and 𝐢𝑝 = 1.548 π‘˜π½/π‘˜π‘”πΎ.Β 


1
Expert's answer
2022-03-15T16:49:23-0400

"A=\frac{\nu RT_1}{\gamma-1}(1-(\frac{p_2}{p_1})^{\gamma-1})=80.6~J."